OpenStudy (anonymous):

if f(x) is greater than 12 but less than 17 ...the int[3,8] f(x) is less than______ and greater than_____? I dont know where to start. please help

OpenStudy (anonymous):

integral from[3,8]

OpenStudy (mathmale):

Suggest you always double-check your typing before you post a problem. As cusemath has just pointed out, there's ambiguity regarding what you mean by "int [3,8]. Because of the [3,8], I interpret "int" to mean INTERVAL here, not integral.

OpenStudy (mathmale):

Here's my first stab at a solution: \[12<f(x)<17\] for all x. Thus, if we restrict ourselves to the interval [3,8], 12<f(x)<17 is still true and is, to my mind at least, the answer you want. Could some genius out there do some thinking and come up with a better solution?

OpenStudy (anonymous):

\[12lef(x)\le17\]

OpenStudy (anonymous):

\[12 \le f(x) \le 17\]

OpenStudy (mathmale):

Cuse: I'm uncertain what you're trying to say. What justifies using "less than or equal to" if the original problem statement was 12<f(x)<17?

OpenStudy (anonymous):

i made a mistake it should be what I just gave

OpenStudy (anonymous):

same for answer. Its not 12 and 17 either

OpenStudy (anonymous):

here it is. 12(8-5) and 17(8-5) are the respective answers I was looking for

OpenStudy (anonymous):

8-3 sorry

OpenStudy (mathmale):

Your providing the answers gives me a bit more insight, but I'm still not entirely sure why these results are the answers. If the max value that function f(x) can have is 17, for any interval, what would be the meaning of 17/(8-3)? At this point I really do not know. Note that 17/(8-3) = 17/5. You typed "f(x)" twice in the problem statement. In the first instance there's no ambiguity. Look at the 2nd appearance of f(x). Is it possible that you left out some wording when typing this problem?